# MAViS: A Multi-Agent Framework for Long-Sequence Video Storytelling

### v4

@@ Abstract
Despite recent advances, long-sequence video generation frameworks still suffer from significant limitations: poor assistive capability, suboptimal visual quality, and limited expressiveness. To mitigate these limitations, we propose MAViS, a multi-agent collaborative framework designed to assist in long-sequence video storytelling by efficiently translating ideas into visual narratives. MAViS orchestrates specialized agents across multiple stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation.In each stage, agents operate under the 3E Principle—Explore, Examine, and Enhance—to ensure the completeness of intermediate outputs. Considering the capability limitations of current generative models, we propose the Script Writing Guidelines to optimize compatibility between scripts and generative tools. Experimental results demonstrate that MAViS achieves state-of-the-art performance in assistive capability, visual quality, and video expressiveness. Its modular framework further enables scalability with diverse generative models and tools. With just a brief idea description, MAViS enables users to rapidly explore diverse visual storytelling and creative directions for sequential video generation by efficiently crafting high-quality, complete long-sequence videos. To the best of our knowledge, MAViS is the only framework that provides multimodal design output – videos with narratives and background music.
@@
Qian Wang Virginia Tech. University USA yqwq1996@vt.edu Ziqi Huang Nanyang Technological University Singapore ZIQI002@e.ntu.edu.sg Ruoxi Jia Virginia Tech. University USA ruoxijia@vt.edug
Paul Debevec Netflix Eyeline Studios USA debevec@scanlinevfx.com Ning Yu††thanks: Corresponding author and project lead Netflix Eyeline Studios USA ning.yu@scanlinevfx.com
@@ 1 Introduction
| Methods | Agentic | Script Writing | Shot Designing | Dynamic | ID Consistency | Audio | Assistive | |
@@
In recent years, video generative models OpenAI (2024); MiniMax (2024); GenmoTeam (2024) have made substantial strides in text-to-video (T2V) and image-to-video (I2V) generation. While these methods demonstrate promising capabilities for shot-level creation, the generation of long videos remains a formidable challenge. Besides, the current scale of computing and data is insufficient to train a single end-to-end model capable of crafting high-quality, minute-long videos.
To overcome this limitation, prior works have explored extending existing video clips KlingAI (2024); Henschel et al. (2024); Ge et al. (2022), or composing sequences of keyframes Liu et al. (2024); Yang et al. (2024b); Rahman et al. (2023); Wang et al. (2024a), or integrating storytelling and video generation Yin et al. (2023b); Zheng et al. (2025). However, these approaches suffer from repetitive motion, lack of narrative cohesion, or result in static image series with limited visual expressions. To better align narrative planning with video animation, MovieDreamer Zhao et al. (2024) adopts an auto-regressive approach to generate key frames from input scripts, while Dreamfactory Xie et al. (2024) and MovieAgent Wu et al. (2025b) propose a hierarchical script–scene–shot pipeline. However, as shown in Table 1, these methods require substantial laborious effort from users for script writing and training LoRA models to maintain ID consistency, which significantly limits their assistive capability and hinders large-scale deployment.
@@
To address this, we introduce MAViS, a multi-agent collaborative framework designed to assist in long-sequence video storytelling by efficiently translating ideas into visual narratives. MAViS encompasses various stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation. Given a brief idea description from the user, the script writing stage first generates a structured script with character definitions and shot outlines. The shot designing stage defines a series of shot-level elements to guide subsequent content generation, while character modeling ensures ID consistency via LoRA-trained appearances. During the keyframe generation stage, T2I models are employed to create the initial static frames for each shot, i.e., the keyframes, which are then expanded into full video sequences in the video animation stage through I2V models, maintaining visual coherence and narrative fluency. Finally, voice-overs and background music (BGM) are added to each shot during the voice generation stage, resulting in a cohesive, long-sequence video.
Crucially, all stages operate under the 3E Principle: - Explore: Initial generation of candidate outputs; - Examine: Reviewing the quality and completeness of the output; - Enhance: Iterative refinement and optimization. After multiple iterations, each stage passes the most complete version of its output to the next, enabling gradual optimization throughout the pipeline.
To mitigate the limitations of current generative models in maintaining background consistency, executing complex actions, and rendering fine-grained details, we propose the Script Writing Guidelines aimed at enhancing the compatibility between scripts and generative models, ultimately improving the expressiveness of the video storytelling. As illustrated in Figure LABEL:fig:teaser, with just a brief idea description, MAViS enables users to rapidly explore diverse visual storytelling and creative directions for sequential video generation by efficiently crafting high-quality, complete long-sequence videos. Furthermore, its modular architecture ensures scalability with various generative tools and models.
@@
We propose MAViS, a multi-agent collaborative framework that encompasses various stages including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation — forming a complete and scalable workflow for long video storytelling.
@@
Experiments demonstrate that MAViS achieves state-of-the-art performance and wins the most user preferences in terms of expressiveness and visual quality. With just a brief idea description from the user, MAViS efficiently generates high-quality and complete long-sequence video storytelling. To the best of our knowledge, MAViS is the only framework that provides multimodal design output.
@@ 2.2 AI Agents
AI agents Yang et al. (2023a); Yin et al. (2023a); Li et al. (2023); Hong et al. (2023); Park et al. (2023), empowered by large language models Brown et al. (2020); Touvron et al. (2023a, b), have shown broad applicability in reasoning, planning, web interaction, and visual generation. In video generation, systems Yang et al. (2024a); Yuan et al. (2024); Wang et al. (2024a) explore world modeling, multi-agent coordination, and retrieval-augmented synthesis, yet often yield static content with limited transitions. Hierarchical pipelines Xie et al. (2024); Wu et al. (2025b) introduce script-to-shot workflows but still rely on laborious scripting and LoRA tuning. In contrast, our MAViS leverages multi-agent collaboration to enable long-sequence video storytelling.
@@ 3 MAViS
We propose MAViS, a multi-agent collaborative framework for long-sequence video storytelling that consists of various stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation.
@@ 3.2 Script Writing Guidelines
Current video generation models struggle to synthesize multiple sequential actions within a single shot (e.g., running followed by walking and then dancing). Therefore, it is advisable to limit each shot to a single, simple action. Furthermore, generative models often fail to craft fine-grained visual elements accurately, such as textual content on smartphone screens or small patterns on objects. Scripts should avoid specifying such details.
@@ 3.3 Script Writing Stage
Upon receiving a brief user prompt, four agents—the Scriptwriter, the Structure Reviewer, the Content Reviewer, and the Style Reviewer—work collaboratively in the script writing stage to form a structured script.
The Scriptwriter first generate a draft script based on the user prompt and the Script Writing Guidelines :
@@ 3.5 Character Modeling Stage
The process begins with the Character Prompt Generator, which generates both text-to-image (T2I) and image-to-video (I2V) prompts for each character. T2I prompts aim to generate a frontal and naturally posed character image, while I2V prompts are used—together with the T2I images—to synthesize character-centric video clips that capture the subject from multiple angles. After both T2I and I2V processes, multi-view character images are sampled from the generated character video. The Caption Generator then generates structured captions for each sampled frame. These images and captions are subsequently used to train LoRA models (if needed).
Both the T2I and I2V processes follow the 3E Principle, but differ from script writing and shot designing in that each iteration generates multiple samples from the generative model pools. Specifically, in the -th iteration, a set of candidate images is generated by the pool of T2I models using the current T2I prompt :
@@
The I2V Evaluator evaluates these candidates according to the video evaluation guide , selects the best one , and provides evaluation result . The I2V Refiner updates the I2V prompt and generates the final video after iterations:
@@ 3.6 Generation Stages
Finally, audio tracks are synchronized with video clips to form the final long-sequence video telling with synchronized visuals and sound: .
@@ 4 Experiment
To evaluate the effectiveness of MAViS, we benchmark it against three three long video generation baselines: VGoT, Mora, and MovieAgent, and conduct detailed ablation studies. We constructed a test set of 20 user prompts, each targeting a one-minute video for evaluation. We employ CLIP, Inception Score, and VBench Huang et al. (2024) for quantitative evaluation, and conduct a user study for subjective assessment. Further experimental settings, metrics, and baseline details are provided in the Appendix.
@@ 4.1.1 Automatic Metrics
Table 2 presents the shot-level evaluation results for both keyframe generation and video animation. MAViS achieves the highest scores in CLIP (4.22) and Inception (12.81) scores, attributed to the enhanced T2I quality enabled by the 3E Principle in the T2I process. In video animation, MAViS slightly lags behind VGoT in subject consistency and background consistency but outperforms all baselines in temporal flickering, motion smoothness, and image quality. This is partly because VGoT crafts videos with limited dynamics and highly saturated visuals, which improves intra-shot consistency and aesthetic quality but at the cost of realism. Overall, MAViS achieves the best comprehensive performance.
@@ 4.2.1 Script Writing Guidelines
Since the evaluators and refiners in both the T2I and I2V processes cannot operate independently, we conduct an ablation study to evaluate the overall impact of the 3E Principle on these modules. The CLIP score, Inception score, Naturalness score (defined as the proportion of generated images or videos that appear physically plausible and naturally rendered), along with Subject Consistency and Dynamic Degree from the VBench metrics, are used to assess key aspects targeted by the T2I and I2V Evaluators—specifically, prompt consistency, visual quality, naturalness, subject consistency, and dynamics. As shown in Table 5, removing the 3E Principle leads to a noticeable drop in performance across both T2I and I2V outputs, particularly in terms of naturalness ( in T2I and in I2V). This highlights the critical role of the 3E Principle in improving generation quality. An exception is observed in the I2V process: the Dynamic Degree is higher without the 3E Principle (). This may be because high-dynamic videos are more prone to physical inconsistencies and instability, leading the agents to favor lower-dynamic, more stable results.
@@ 4.2.3 Shot Reviewer, Voice Reviewer, and Subtitle Refiner
5 Conclution
In this work, we presented MAViS, a multi-agent collaborative framework designed to assist in long-sequence video storytelling. By orchestrating multiple stages within a unified pipeline, MAViS effectively addresses the key limitations of prior approaches — poor assistive capability, limited visual expressiveness, and incomplete outputs. Central to our framework is the 3E Principle (Explore, Examine, Enhance), which enables iterative refinement across all stages, maximizing output quality. Additionally, our proposed Script Writing Guidelines improve compatibility between narrative structure and generative tools. Experimental results validate that MAViS significantly advances the assistive capability, visual quality, and expressiveness of long-sequence video storytelling.
@@ A.3 AI Agents
AI agents, empowered by large language models Brown et al. (2020); Touvron et al. (2023a, b), have shown broad applicability in reasoning, planning, web interaction, and visual generation. General-purpose LLMs underpin agent frameworks Yang et al. (2023a); Yin et al. (2023a); Li et al. (2023); Hong et al. (2023); Park et al. (2023), while multimodal models Achiam et al. (2023); Anil et al. (2023) extend agent capabilities across modalities Yang et al. (2023b); Zhang and Zhang (2024); Zheng et al. (2024); Wang et al. (2024e); Liu et al. (2025); Wang et al. (2024d); Hang et al. (2024); Liao et al. (2025). In video generation, systems Yang et al. (2024a); Yuan et al. (2024); Wang et al. (2024a) explore world modeling, multi-agent coordination, and retrieval-augmented synthesis, yet often yield static content with limited transitions. Hierarchical pipelines Xie et al. (2024); Wu et al. (2025b) introduce script-to-shot workflows but still rely on laborious scripting and LoRA tuning.
@@ Appendix B Explanation and Discuss about the Script Writing Guideline
For more details, refer to the Script Writing Guideline, please refer to the system messages.
To evaluate the performance of MAViS, we compared it with three open-sourced long video generation frameworks. Further experimental details and results are provided in the Appendix.

### v5

@@ Abstract
Despite recent advances, long-sequence video generation frameworks still suffer from significant limitations: poor assistive capability, suboptimal visual quality, and limited expressiveness. To mitigate these limitations, we propose MAViS, an end-to-end multi-agent collaborative framework for long-sequence video storytelling. MAViS orchestrates specialized agents across multiple stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation.In each stage, agents operate under the 3E Principle—Explore, Examine, and Enhance—to ensure the completeness of intermediate outputs. Considering the capability limitations of current generative models, we propose the Script Writing Guidelines to optimize compatibility between scripts and generative tools. Experimental results demonstrate that MAViS achieves state-of-the-art performance in assistive capability, visual quality, and video expressiveness. Its modular framework further enables scalability with diverse generative models and tools. With just a brief prompt, MAViS enables users to rapidly explore diverse visual storytelling and creative directions for sequential video generation by efficiently producing high-quality, complete long-sequence videos. To the best of our knowledge, MAViS is the only framework that provides multimodal design output—videos with narratives and background music.
@@
Qian Wang Virginia Tech USA yqwq1996@vt.edu Ziqi Huang Nanyang Technological University Singapore ZIQI002@e.ntu.edu.sg Ruoxi Jia Virginia Tech USA ruoxijia@vt.edu
Paul Debevec Eyeline Studios USA debevec@scanlinevfx.com Ning Yu Eyeline Studios USA ning.yu@scanlinevfx.com
@@ 1 Introduction
| Methods | Agentic | Script Writing | Shot Designing | Dynamic | ID Consistency | Audio | Automated | |
@@
In recent years, video generative models OpenAI (2024); MiniMax (2024); GenmoTeam (2024) have made substantial strides in text-to-video (T2V) and image-to-video (I2V) generation. While these methods demonstrate promising capabilities for shot-level creation, the generation of long videos remains a formidable challenge. Besides, the current scale of computing and data is insufficient to train a single end-to-end model capable of producing high-quality, minute-long videos.
To overcome this limitation, prior works have explored extending existing video clips KlingAI (2024); Henschel et al. (2024); Ge et al. (2022), or composing sequences of keyframes Liu et al. (2024); Yang et al. (2024b); Rahman et al. (2023); Wang et al. (2024a), or integrating storytelling and video generation Yin et al. (2023b); Zheng et al. (2025). However, these approaches suffer from repetitive motion, lack of narrative cohesion, or result in static image series with limited visual expressions. To better align narrative planning with video animation, MovieDreamer Zhao et al. (2024) adopts an auto-regressive approach to generate key frames from input scripts, while Dreamfactory Xie et al. (2024) and MovieAgent Wu et al. (2025b) propose a hierarchical script–scene–shot pipeline. However, as shown in Table 1, these methods require substantial manual effort from users for script writing and training LoRA models to maintain ID consistency, which significantly limits their assistive capability and hinders large-scale deployment.
@@
To address this, we introduce MAViS, a multi-agent framework for end-to-end long-sequence video generation. MAViS encompasses various stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation. Given a brief user prompt, the script writing stage first generates a structured script with character definitions and shot outlines. The shot designing stage defines a series of shot-level elements to guide subsequent content generation, while character modeling ensures ID consistency via LoRA-trained appearances. During the keyframe generation stage, T2I models are employed to create the initial static frames for each shot, i.e., the keyframes, which are then expanded into full video sequences in the video animation stage through I2V models, maintaining visual coherence and narrative fluency. Finally, voice-overs and background music (BGM) are added to each shot during the voice generation stage, resulting in a cohesive, long-sequence video.
Crucially, all stages operate under the 3E Principle: - Explore: Initial generation of candidate outputs; - Examine: Reviewing the quality and completeness of the output; - Enhance: Iterative refinement and optimization. After multiple iterations, each stage passes the most complete version of its output to the next, enabling gradual optimization throughout the end-to-end pipeline.
To mitigate the limitations of current generative models in maintaining background consistency, executing complex actions, and rendering fine-grained details, we propose the Script Writing Guidelines aimed at enhancing the compatibility between scripts and generative models, ultimately improving the expressiveness of the video storytelling. As illustrated in Figure LABEL:fig:teaser, with just a brief prompt, MAViS enables users to rapidly explore diverse visual storytelling and creative directions for sequential video generation by efficiently producing high-quality, complete long-sequence videos. Furthermore, its modular architecture ensures scalability with various generative tools and models.
@@
We propose MAViS, an end-to-end multi-agent framework that encompasses various stages including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation — forming a complete and scalable workflow for long video storytelling.
@@
Experiments demonstrate that MAViS achieves state-of-the-art performance and receive the most user preference votes in terms of expressiveness and visual quality. With just a brief user prompt, MAViS efficiently generates high-quality and complete long-sequence video storytelling. To the best of our knowledge, MAViS is the only framework that provides multimodal design output.
@@ 2.2 AI Agents
AI agents Yang et al. (2023a); Yin et al. (2023a); Li et al. (2023); Hong et al. (2023); Park et al. (2023), empowered by large language models Brown et al. (2020); Touvron et al. (2023a, b), have shown broad applicability in reasoning, planning, web interaction, and visual generation. In video generation, systems Yang et al. (2024a); Yuan et al. (2024); Wang et al. (2024a) explore world modeling, multi-agent coordination, and retrieval-augmented synthesis, yet often yield static content with limited transitions. Hierarchical pipelines Xie et al. (2024); Wu et al. (2025b) introduce script-to-shot workflows but still rely on manual scripting and LoRA tuning. In contrast, our MAViS leverages multi-agent collaboration to enable long-sequence video storytelling.
@@ 3 MAViS
We propose MAViS, a multi-agent end-to-end framework for long-sequence video storytelling that consists of various stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation.
@@ 3.2 Script Writing Guidelines
Current video generation models struggle to synthesize multiple sequential actions within a single shot (e.g., running followed by walking and then dancing). Therefore, it is advisable to limit each shot to a single, simple action. Furthermore, generative models often fail to produce fine-grained visual elements accurately, such as textual content on smartphone screens or small patterns on objects. Scripts should avoid specifying such details.
@@ 3.3 Script Writing Stage
Upon receiving a brief user prompt, four agents—the Scriptwriter, the Structure Reviewer, the Content Reviewer, and the Style Reviewer—work collaboratively in the script writing stage to produce a structured script.
The Scriptwriter first generates a draft script based on the user prompt and the Script Writing Guidelines :
@@ 3.5 Character Modeling Stage
The process begins with the Character Prompt Generator, which generates both text-to-image (T2I) and image-to-video (I2V) prompts for each character. T2I prompts aim to generate a frontal and naturally posed character image, while I2V prompts are used—together with the T2I images—to synthesize character-centric video clips that capture the subject from multiple angles. After both T2I and I2V processes, multi-view character images are sampled from the generated character video. The Caption Generator then produces structured captions for each sampled frame. These images and captions are subsequently used to train LoRA models (if needed).
Both the T2I and I2V processes follow the 3E Principle, but differ from script writing and shot designing in that each iteration produces multiple samples from the generative model pools. Specifically, in the -th iteration, a set of candidate images is generated by the pool of T2I models using the current T2I prompt :
@@
The I2V Evaluator evaluates these candidates according to the video evaluation guide , selects the best one , and provides evaluation result . The I2V Refiner updates the I2V prompt and produce the final video after iterations:
@@ 3.6 Generation Stages
Finally, audio tracks are synchronized with video clips to produce the final long-sequence video telling with synchronized visuals and sound: .
@@ 4 Experiment
To evaluate the effectiveness of MAViS, we benchmark it against three long video generation baselines: VGoT, Mora, and MovieAgent, and conduct detailed ablation studies. We constructed a test set of 20 user prompts, each targeting a one-minute video for evaluation. We employ CLIP, Inception Score, and VBench Huang et al. (2024) for quantitative evaluation, and conduct a user study for subjective assessment. Further experimental settings, metrics, and baseline details are provided in the Appendix.
@@ 4.1.1 Automatic Metrics
Table 2 presents the shot-level evaluation results for both keyframe generation and video animation. MAViS achieves the highest scores in CLIP (34.22) and Inception (12.81) scores, attributed to the enhanced T2I quality enabled by the 3E Principle in the T2I process. In video animation, MAViS slightly lags behind VGoT in subject consistency and background consistency but outperforms all baselines in temporal flickering, motion smoothness, and image quality. This is partly because VGoT produces videos with limited dynamics and highly saturated visuals, which improves intra-shot consistency and aesthetic quality but at the cost of realism. Overall, MAViS achieves the best comprehensive performance.
@@ 4.2.1 Script Writing Guidelines
Since the evaluators and refiners in both the T2I and I2V processes cannot operate independently, we conduct an ablation study to evaluate the overall impact of the 3E Principle on these modules. The CLIP score, Inception score, Naturalness score (defined as the proportion of generated images or videos that appear physically plausible and naturally rendered), along with Subject Consistency and Dynamic Degree from the VBench metrics, are used to assess key aspects targeted by the T2I and I2V Evaluators. As shown in Table 5, removing the 3E Principle leads to a noticeable drop in performance across both T2I and I2V outputs, particularly in terms of naturalness ( in T2I and in I2V). This highlights the critical role of the 3E Principle in improving generation quality. An exception is observed in the I2V process: the Dynamic Degree is higher without the 3E Principle (). This may be because high-dynamic videos are more prone to physical inconsistencies and instability, leading the agents to favor lower-dynamic, more stable results.
@@ 4.2.3 Shot Reviewer, Voice Reviewer, and Subtitle Refiner
5 Conclusion
In this work, we presented MAViS, a multi-agent end-to-end framework for long-sequence video storytelling. By orchestrating multiple stages within a unified pipeline, MAViS effectively addresses the key limitations of prior approaches — poor assistive capability, limited visual expressiveness, and incomplete outputs. Central to our framework is the 3E Principle (Explore, Examine, Enhance), which enables iterative refinement across all stages, maximizing output quality. Additionally, our proposed Script Writing Guidelines improve compatibility between narrative structure and generative tools. Experimental results validate that MAViS significantly advances the assistive capability, visual quality, and expressiveness of long-sequence video storytelling.
Limitations
In practical AI short film production, there are multiple ways to generate a shot—for example, using direct T2V, combining several I2V generations from different camera angles, or applying image editing tools and image generation models with stronger semantic control and compositional reasoning (e.g., nanobanana Google (2025) and HunyuanImage 3.0 Cao et al. (2025)) to enforce identity and background consistency. However, this paper restricts long-sequence video storytelling to a T2I+I2V paradigm, and the scriptwriting guidelines are explicitly adapted to accommodate this pipeline. While this design simplifies the workflow, it also limits shot diversity, thereby constraining the expressive capacity of the generated storytelling.
Moreover, current storylines lack interactions and dialogue between characters, and the absence of video editing mechanisms in the framework further degrades the cinematic quality of the final output.
To address these limitations, future work will allow agents to autonomously select appropriate generation strategies and tools, and will incorporate character dialogues, scene-specific sound effects, and video editing functionalities, aiming to produce more expressive and coherent video storytelling.
Another limitation of this work lies in the limited reliability of the T2I and I2V Evaluators. Current multimodal foundation models still exhibit limited accuracy in assessing visual content, resulting in suboptimal scoring performance when ranking candidate outputs. To mitigate this in practice, we incorporate additional evaluation operators—such as those from VBench for assessing prompt consistency and image quality—to assist the Evaluators. However, to ensure fairness during experimental evaluation, these operators were not used. In future work, we plan to integrate a broader set of evaluation operators to enhance the overall assessment process.
Ethical Considerations
While multi-agent frameworks like MAViS enable efficient and creative video storytelling, AI-generated media must be used responsibly. Because MAViS integrates multiple open-source models and commercial APIs, its outputs may inherit safety risks, biases, or limitations from these components. We encourage transparent disclosure of model provenance, dataset sources, and responsible use practices when employing MAViS in creative or research contexts.
Potential Risks.
MAViS is designed as a research framework for controllable video generation and multimodal reasoning. However, its capacity to produce coherent, realistic long-sequence videos may be misused to create misleading or harmful synthetic content, such as disinformation, deepfakes, or fabricated narratives. In addition, because it relies on large pretrained models and APIs, MAViS may inadvertently propagate existing societal biases, cultural stereotypes, or skewed representations into generated stories. We emphasize that MAViS is intended solely for academic research and creative exploration under ethical guidelines, not for deployment in sensitive, deceptive, or adversarial applications.
@@ A.3 AI Agents
AI agents, empowered by large language models Brown et al. (2020); Touvron et al. (2023a, b), have shown broad applicability in reasoning, planning, web interaction, and visual generation. General-purpose LLMs underpin agent frameworks Yang et al. (2023a); Yin et al. (2023a); Li et al. (2023); Hong et al. (2023); Park et al. (2023), while multimodal models Achiam et al. (2023); Anil et al. (2023) extend agent capabilities across modalities Yang et al. (2023b); Zhang and Zhang (2024); Zheng et al. (2024); Wang et al. (2024e); Liu et al. (2025); Wang et al. (2024d); Hang et al. (2024); Liao et al. (2025). In video generation, systems Yang et al. (2024a); Yuan et al. (2024); Wang et al. (2024a) explore world modeling, multi-agent coordination, and retrieval-augmented synthesis, yet often yield static content with limited transitions. Hierarchical pipelines Xie et al. (2024); Wu et al. (2025b) introduce script-to-shot workflows but still rely on manual scripting and LoRA tuning.
@@ Appendix B Explanation and Discuss about the Script Writing Guideline
| Stage | Agent/Model | Role | Reviewer | Refiner |
| Script Writing | Scriptwriter | Write script | Structure Reviewer, | |
| Content Reviewer, | - | |||
| Style Reviewer | ||||
| Shot Designing | Shot Designer | Design each shot of the script | Shot Reviewer | - |
| Keyframe Generation | T2I Prompt Generator | Generate T2I prompt for each shot | - | - |
| Keyframe Generation, | T2I Models | Text-to-image generation for each shot/character | T2I Evaluator | T2I Refiner |
| Character Modeling | ||||
| Video Animation, | I2V Prompt Generator | Generate I2V prompt for each shot/character | - | - |
| Character Modeling | ||||
| Video Animation | I2V Models | Image-to-video generation for each shot | I2V Evaluator | I2V Refiner |
| Character Modeling | Character Prompt Generator | Generate T2I and I2V prompts for each character | - | - |
| Character Modeling | Caption Generator | Generate captions for each character images | - | - |
| Audio Generation | Voice Designer | Choose voice ID, emotion, and BGM | Voice Reviewer | - |
| Audio Generation | T2A Model | Text-to-audio generation for each shot | - | Subtitle Refiner |